OpenStudy (naga447):

graphing help!!!!! I'm trying to find two points on a graph that the y intersect is equal to -2 can anyone help? (this is all the information that they gave me)

OpenStudy (anonymous):

Put it into slope-intercept form. (y=mx+b)

OpenStudy (naga447):

-2 = mx + b is this correct?

OpenStudy (anonymous):

Is this multiple choice or?

OpenStudy (anonymous):

No, the slope intercept form would be y=mx-2.

OpenStudy (anonymous):

Now, do they give you points to use, like is this multiple choice?

OpenStudy (naga447):

no its not multiple choice this is everything there is to the question...

OpenStudy (anonymous):

That means you could create your own slope.

OpenStudy (anonymous):

Now, you would graph it, and just choose 2 points in intercepts, and those would be 2 of your points.

OpenStudy (naga447):

Okay that makes sense thank you!

OpenStudy (anonymous):

Mhm, no problem ;)

OpenStudy (masumanwar):

write like x/a+y/b=1

OpenStudy (naga447):

If I made the slope equal to 0 would 2, -2 and 0, -2 work?

OpenStudy (anonymous):

If the slope it zero, that means y = -2, which means it will be undefined.

OpenStudy (anonymous):

That means the graph would be a straight horizontal line going through -2, which means yes, yours would work

OpenStudy (naga447):

Okay thank you for checking them!!!

OpenStudy (anonymous):

Np ^_~